### v2

# Why gauge invariance applies to statistical mechanics

Abstract
We give an introductory account of the recently identified gauge invariance of the equilibrium statistical mechanics of classical many-body systems [J. Müller et al.arXiv:2406.19235, Phys. Rev. Lett. (to appear)]. The gauge transformation is a non-commutative shifting operation on phase space that keeps the differential phase space volume element and hence the Gibbs integration measure conserved. When thermally averaged any observable is an invariant, including thermodynamic and structural quantities. Shifting transformations are canonical in the sense of classical mechanics. They also form an infinite-dimensional group with generators of infinitesimal transformations that build a non-commutative Lie algebra. We lay out the connections with the underlying geometry of coordinate displacement and with Noether’s theorem. Spatial localization of the shifting yields differential operators that satisfy commutator relationships, which we describe both in purely configurational and in full phase space setups. Standard operator calculus yields corresponding equilibrium hyperforce correlation sum rules for general observables and order parameters. Using Monte Carlos simulations we demonstrate explicitly the gauge invariance for finite shifting. We argue in favour of using the gauge invariance as a statistical mechanical construction principle for obtaining exact results and for formulating smart sampling algorithms.
I Introduction
Gauge field theories are fundamental for our understanding of nature. Considering the invariance against well-defined gauge transformations constitutes a universal construction principle, which formalizes the independence of the physical predictions that a theory provides from the precise choice of the variables that are in use. The perhaps most well-known case of gauge invariance affects the (scalar and vector) potentials of electrodynamics. While thereby specific choices of gauge can significantly simplify practical calculations, arguably as important are the fundamental consequences of gauge invariance. In the present case it is the intimate connection with the local conservation of electrical charge. We describe this standard case below as a template for the many-body physics that we address in this contribution.
Noether’s theorem of invariant variations noether1918 ; byers1998 ; brading2002 ; read2022book provides the appropriate mathematical framework for systematically addressing the consequences that follow from the juxtaposition of inherent independence and apparent dependence on the choice of gauge. While forming a staple of field theories for many decades, Noether’s theorem noether1918 ; byers1998 ; brading2002 ; read2022book has seen only relatively recently an increasing number of applications to statistical physics revzen1970 ; baez2013markov ; marvian2014quantum ; sasa2016 ; sasa2019 ; bravetti2023 ; budkov2022 ; brandyshev2023 ; budkov2024ionic ; budkov2024jcp ; beyen2024 ; beyen2024generic , both in and out-of-equilibium. As much of statistical physics is based on insightful combination of suitable approximations with exact results, perhaps most notably in the form of equilibrium sum rules hansen2013 ; evans1979 ; evans1992 ; baus1984 ; evans1990 ; henderson1992 ; triezenberg1972 ; percus1962 ; gul2024testParticle , there is arguably much potential for making progress.
Applications of the Noether theorem were performed in a variety of statistical mechanical settings as follows. In an early study of functional integrals in statistical physics and primarily based on quantum field theory, Revzen derived the continuity equation from Noether’s theorem revzen1970 . Baez and Fong formulated a Noether theorem for Markov processes baez2013markov . Marvian and Spekkens provided an extension of Noether’s theorem by quantifying the asymmetry of quantum states marvian2014quantum . Sasa and Yokokura formulated the thermodynamic entropy as a Noether invariant sasa2016 . Their work lead to investigation of the thermodynamical path integral and emergent symmetry sasa2019 . Recently the thermodynamic entropy was viewed as a Noether invariant from contact geometry bravetti2023 . Budkov and collaborators carried out a range of studies of complex systems, where the use of Noether’s theorem was crucial. This work addressed modified Poisson-Boltzmann equations and macroscopic forces in inhomogeneous ionic fluids budkov2022 , as well as a covariant field theory of mechanical stresses in inhomogeneous ionic fluids brandyshev2023 . The thermomechanical approach was applied to calculating mechanical stresses in inhomogeneous ionic fluids budkov2024ionic and the surface tension of aqueous electrolyte solutions budkov2024jcp . In recent work, Beyen and Maes identified the entropy as a Noether charge for quasistatic gradient flow beyen2024 and for more general setups beyen2024generic .
The statistical mechanics of particle-based systems is well accessible via direct simulation frenkel2023book . Thereby inherent statistical uncertainties can be reduced considerably by a range of recent methods, most prominently via mapped averaging moustafa2015 ; schultz2016 ; moustafa2017jctp ; moustafa2017prb ; schultz2018 ; purohit2018 ; moustafa2019 ; purohit2020 ; moustafa2022 ; lin2018 ; trokhymchuk2019 ; schultz2019 ; purohit2019 and force sampling frenkel2023book ; borgis2013 ; delasheras2018forceSampling ; coles2019 ; coles2021 ; rotenberg2020 ; mangaud2020 ; coles2023revelsMD ; renner2023torqueSampling . Mapped averaging, as pioneered by Kofke and coworkers moustafa2015 ; schultz2016 ; moustafa2017jctp ; moustafa2017prb ; schultz2018 ; purohit2018 ; moustafa2019 ; purohit2020 ; moustafa2022 ; lin2018 ; trokhymchuk2019 ; schultz2019 ; purohit2019 , is a systematic means for the reformulation of ensemble averages via flexible (coordinate) mappings schultz2016 . The approach was used in a variety of contexts, including sampling of thermal properties of crystals moustafa2015 ; schultz2016 ; moustafa2017jctp ; moustafa2017prb ; schultz2018 ; purohit2018 ; moustafa2019 ; purohit2020 ; moustafa2022 , of liquid water lin2018 , and of hard sphere trokhymchuk2019 and Lennard-Jones fluids schultz2018 . Force sampling originates from a pioneering early investigation by Borgis et al. borgis2013 and we refer to the review by Rotenberg rotenberg2020 for a comprehensive account of the approach. As a central concept, force sampling is based on using numerical integration methods to invert spatial gradients of correlation functions with the aim to exploit the accompanying smoothening effect. The relationship of force-sampling methods and mapped averaging has been illuminated in Ref. purohit2019 .
A specific shifting operation on the phase space associated with general interacting many-body systems was recently put forward and shown to be useful in a broad range of statistical mechanical settings hermann2021noether ; hermann2022topicalReview ; hermann2022variance ; hermann2022quantum ; sammueller2023whatIsLiquid ; hermann2023whatIsLiquid ; robitschko2024any . These applications include formulating force-based density functional theory tschopp2022forceDFT ; sammueller2022forceDFT , self-consistency conditions for neural functionals sammueller2023neural ; sammueller2023whyNeural , the formulation of hyperdensity functional theory for the description of the equilibrium behaviour of general observables sammueller2024hyperDFT ; sammueller2024whyhyperDFT , as well as nonequilibrium power functional schmidt2022rmp sum rules hermann2021noether ; delasheras2023perspective ; zimmermann2024ml .
The phase space shifting hermann2021noether ; hermann2022topicalReview ; hermann2022variance ; hermann2022quantum ; sammueller2023whatIsLiquid ; hermann2023whatIsLiquid ; robitschko2024any ; tschopp2022forceDFT ; sammueller2022forceDFT was recently argued to constitute a local gauge transformation of statistical mechanical microstates mueller2024gauge . Here we provide in-depth background for this gauge shifting transformation mueller2024gauge . The transformation is geometric in nature and we describe in detail its origins in spatial displacements of particle coordinates. We introduce a purely configurational version and show that including the complementary momentum transformation mueller2024gauge , realized by (position-resolved) matrix multiplication, completes a very specific mathematical structure. As its key features, the shifting transformation is both canonical in the sense of classical mechanics and it constitutes a non-commutative continuous group. Phase space shifting leaves any equilibrium average invariant, which establishes its status as a gauge transformation. Meeting expectations for this role, the shifting unambiguously leads to a breadth of exact identities.
Our gauge theory carries strong similarities in its mathematical structure with the theory of Lie groups and corresponding Lie algebras of infinitesimal transformations. Important examples of nontrivial Lie groups in physics include the rotation matrices of three-dimensional space and the Pauli matrices to describe quantum mechanical spin 1/2 particles goldstein2002 . Our theory goes in important ways further than these elementary cases. In particular the tight interweaving with spatial Dirac distributions is notable. This feature provides microscopically sharp resolution and, as we demonstrate, it allows one to retain particle-level precision for all ensuing force and hyperforce correlation functions and their sum rules.
The paper is organized as follows. We describe in Sec. II the standard gauge invariance of classical electrodynamics to both provide a discussion of general concepts of gauge invariance and also to formulate a blueprint for the subsequent particle-based physics. Readers who are familiar with this material are welcome to skip forward to Sec. III where we describe the statistical mechanical phase space gauge invariance. The underlying phase space shifting transform being presented in Sec. III.1. Its central properties are derived in Sec. III.2 and the transform is thereby taken to be in general of finite magnitude. Rich additional mathematical structure is revealed when considering infinitesimal transforms as is described in detail for particle configuration (position) shifting in Sec. III.3.
We turn to statistical mechanics in Sec. IV and first define the general concepts, including the partition sum and thermal averages in Sec. IV.1. The relevant standard one-body observables are introduced in Sec. IV.2. Configuration shifting is applied to derive exact sum rules in Sec. IV.3. General phase space shifting, which includes additional momentum displacement and restores the full canonical phase space transform, is described in Sec. IV.4. We present implications for and applications to computer simulation work in Sec. V. We give our conclusions and an outlook on possible future work in Sec. VI.
Figure 1: Group structure of the gauge transformations in electrodynamics (curves) and in statistical mechanics (arrows), depicted for global transforms for the latter case. Shown are different forms of the scalar gauge function $\varphi(x)$ that parameterizes the transform in electrodynamics. A given gauge function $\varphi_{1}(x)$ has an inverse $-\varphi_{1}(x)$. Two transforms $\varphi_{1}(x)$ and $\varphi_{2}(x)$ combine additively to $\varphi_{21}(x)=\varphi_{1}(x)+\varphi_{2}(x)$. For the statistical mechanics the global gauge invariance is shown for different realizations of a constant shifting vector ${\boldsymbol{\epsilon}}_{1}$. The inverse transform ${\boldsymbol{\epsilon}}_{-1}$ is represented by the negative shifting vector, ${\boldsymbol{\epsilon}}_{-1}=-{\boldsymbol{\epsilon}}_{1}$. Two transformations that correspond to ${\boldsymbol{\epsilon}}_{1}$ and ${\boldsymbol{\epsilon}}_{2}$ combine additively to ${\boldsymbol{\epsilon}}_{21}={\boldsymbol{\epsilon}}_{1}+{\boldsymbol{\epsilon}}_{2}$. Also illustrated is the conserved phase space volume element $d{\mathbf{r}}_{i}d{\bf p}_{i}$.
II Classical electrodynamics
The prototypical and arguably most well-known example of gauge invariance in physics is that of classical electrodynamics kobe1980 ; brading2002 ; iqbal2024 , where the position- and time-dependent scalar electric potential $V({\mathbf{r}},t)$ and the magnetic vector potential ${\bf A}({\mathbf{r}},t)$ are transformed on the basis of a scalar gauge function $\varphi({\mathbf{r}},t)$. The potentials are transformed according to the map:
$V({\mathbf{r}},t)$ $\to V({\mathbf{r}},t)-\partial\varphi({\mathbf{r}},t)/\partial t,$ (1)
${\bf A}({\mathbf{r}},t)$ $\to{\bf A}({\mathbf{r}},t)+\nabla\varphi({\mathbf{r}},t),$ (2)
where $\nabla$ indicates the derivative with respect to position ${\mathbf{r}}$.
The form of the gauge function $\varphi({\mathbf{r}},t)$ can be arbitrarily chosen, as it has no discernible effect on Maxwell’s equations, which govern the dynamics of the magnetic induction ${\bf B}({\mathbf{r}},t)=\nabla\times{\bf A}({\mathbf{r}},t)$ and of the electrical field ${\bf E}({\mathbf{r}},t)=-\nabla V({\mathbf{r}},t)-\partial{\bf A}({\mathbf{r}},t)/\partial t$. This gauge invariance can readily be verified. That the transformation preserves the form of ${\bf B}({\mathbf{r}},t)$ follows from elementary vector calculus via ×$\nabla\times\nabla\varphi({\mathbf{r}},t)=0$. Likewise ${\bf E}({\mathbf{r}},t)$ remains unchanged as the mixed partial derivative of $\varphi({\mathbf{r}},t)$ with respect to ${\mathbf{r}}$ and $t$ can be interchanged and hence the two terms that emerge from the gauge transformation cancel each other such that $\nabla\partial\varphi({\mathbf{r}},t)/\partial t-\partial\nabla\varphi({\mathbf{r}},t)/\partial t=0$.
The gauge freedom of choice of the form of $\varphi({\mathbf{r}},t)$ has very profound implications quantum mechanically and a rich structure of modern theoretical physics is based on different forms of gauge symmetry. As background here we lay out the gauge freedom for the classical electrodynamical case; Ref. stokes2022 reviews gauge freedom in quantum electrodynamics.
We use the standard covariant formulation with the four-vector potential $A_{\nu}(x)=(V({\mathbf{r}},t)/c,{\bf A}({\mathbf{r}},t))$ where the four-vector $x_{\nu}=(ct,{\mathbf{r}})$ indicates the spacetime point, $c$ is the speed of light, and $\nu$ denotes the component of spacetime. The electrical charge density $q({\mathbf{r}},t)$ and the electrical current density ${\bf J}({\mathbf{r}},t)$ form a four-current $J_{\nu}(x)=(cq({\mathbf{r}},t),{\bf J}({\mathbf{r}},t))$. The electromagenetic field tensor then follows as $F_{\nu\lambda}(x)=\partial_{\nu}A_{\lambda}(x)-\partial_{\lambda}A_{\nu}(x)$, where $\partial_{\nu}$ indicates the derivative with respect to $x^{\nu}$, which in explicit form is $\partial_{\nu}=(c^{-1}\partial/\partial t,-\nabla)$. All Greek indices range from 0 (time-like) to 1, 2, and 3 (all space-like) and the metric has signature (+,,,)(+,-,-,-)( + , - , - , - ).
This covariant formalism allows one to express the gauge transformation equations (1) and (2) in the succinct form
$A_{\nu}(x)\to A_{\nu}(x)+\partial_{\nu}\varphi(x).$ (3)
It is revealing to address the consequences of Eq. (3) on the standard electromagnetic action $S$ which consists of contributions that arise from free fields ($S_{\rm free}$) and external influence ($S_{\rm ext}$) according to the sum $S=S_{\rm free}+S_{\rm ext}$. The two individual terms are given by
[…]
where $\mu_{0}$ is the magnetic permeability of the vacuum and we use Einstein summation convention over pairs of Greek spacetime indices. Requiring stationarity of $S$ against changes in $A_{\nu}(x)$ yields upon multiplying by $\mu_{0}$ the Maxwell equations in the compact covariant form
$\partial_{\nu}F^{\nu\lambda}(x)$ $=\mu_{0}J^{\lambda}(x).$ (6)
The free field action $S_{\rm free}$ is already an invariant and we hence investigate the effects of the gauge transformation on the external action $S_{\rm ext}$. We apply the transform (3) to Eq. (5) which yields
[…]
In Eq. (7) we have identified the first term in the integral as $S_{\rm ext}$ via Eq. (5) and in the second term we have integrated by parts and assumed that boundary terms vanish.
Imposing that $S_{\rm ext}$ shall be gauge invariant implies that the second term on the right hand side of Eq. (8) must vanish. Given that the form of the gauge function $\varphi(x)$ is arbitrary, this can only be true provided that
$\partial_{\nu}J^{\nu}(x)$ $=0,$ (9)
which is the continuity equation for the charge distribution. When returning to three-vectors Eq. (9) attains the familiar more elementary form $\partial q({\mathbf{r}},t)/\partial t=-\nabla\cdot{\bf J}({\mathbf{r}},t)$.
The simplicity of the specific form (5) of the external action allowed us to derive Eq. (9) directly. Instead of relying explicitly on the fundamental lemma of variational calculus, we can equivalently utilize functional differentiation, here with respect to the gauge function $\varphi(x)$. We lay out these connections in appendix A, where we also demonstrate the gauge invariance of $S_{\rm free}$ explicitly.
Gauge transformations encompass specific group structure that is associated with repeated application of the transformation. In particular, chaining together two subsequent transforms constitutes again a single transform. This is easy to see when considering two transforms represented by two different gauge functions $\varphi_{1}(x)$ and $\varphi_{2}(x)$. Then the composite transform is simply characterized by the sum $\varphi_{21}(x)=\varphi_{1}(x)+\varphi_{2}(x)$. Clearly, the order of the first and the second transformations is irrelevant, $\varphi_{21}(x)=\varphi_{12}(x)=\varphi_{2}(x)+\varphi_{1}(x)$, and hence the group is commutative (Abelian). For completeness, a transform with given $\varphi_{1}(x)$ has an inverse transform characterized by $-\varphi_{1}(x)$, the neutral group element is $\varphi(x)=0$, and the composition is associative. These properties formally establish the mathematical group structure.
An illustration of the local gauge transformation of the potentials of electrodynamics is shown in Fig. 1. We also show a depiction of global shifts that displace the coordinates ${\mathbf{r}}_{i}$ of particle $i$ in a many-body system. Statistical mechanical averages were shown to be invariant under such transformations of all particles hermann2021noether ; hermann2022topicalReview ; hermann2022variance and we refer the reader to Refs. brandyshev2023 ; brandyshev2024 for similar applications in field theory. In the following we turn to the mathematically much richer local version of phase space shifting hermann2022quantum ; sammueller2023whatIsLiquid ; hermann2023whatIsLiquid ; robitschko2024any ; tschopp2022forceDFT ; sammueller2022forceDFT ; mueller2024gauge .
III Gauge invariance of phase space
III.1 Local shifting as a canonical transformation
As laid out in the introduction, Noether’s theorem noether1918 ; byers1998 ; brading2002 ; read2022book has been applied in statistical physics in a variety of settings revzen1970 ; baez2013markov ; marvian2014quantum ; sasa2016 ; sasa2019 ; bravetti2023 ; budkov2022 ; brandyshev2023 ; budkov2024ionic ; budkov2024jcp ; beyen2024 ; beyen2024generic . Here we describe the phase space shifting transformation hermann2021noether ; hermann2022topicalReview ; hermann2022variance ; tschopp2022forceDFT ; hermann2022quantum ; sammueller2023whatIsLiquid ; hermann2023whatIsLiquid ; robitschko2024any that affects the positions ${\mathbf{r}}_{i}$ and the momenta ${\bf p}_{i}$ of a classical many-body system, where the index $i=1,\ldots,N$ enumerates the $N$ particles. The shifting transform hermann2022quantum is parameterized by a smooth (infinitely differentiable) three-dimensional vector field ${\boldsymbol{\epsilon}}({\mathbf{r}})$, where ${\mathbf{r}}$ denotes spatial position.
The position and momentum degrees of freedom of each particle $i$ are affected in the same way. Specifically, the phase space transformation is
${\mathbf{r}}_{i}$ $\to{\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})={\mathbf{r}}_{i}^{\prime},$ (10)
${\bf p}_{i}$ $\to[{\mathbbm{1}}+\nabla_{i}{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})]^{-1}\cdot{\bf p}_{i}={\bf p}_{i}^{\prime},$ (11)
where ${\mathbbm{1}}$ denotes the $d\times d$-unit matrix, with $d$ indicating spatial dimensionality, $\nabla_{i}$ denotes the derivative with respect to particle position ${\mathbf{r}}_{i}$, the superscript $-1$ denotes matrix inversion, and the transformed variables are indicated by a prime (rather than the tilde used in Refs. mueller2024gauge ; hermann2023whatIsLiquid ; robitschko2024any ). We adopt the convention that the $ab$-component of $\nabla_{i}{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})$ is $[\nabla_{i}{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})]_{ab}=\nabla_{i,a}\epsilon_{b}({\mathbf{r}}_{i})$, where the indices $a$,$b$ denote the Cartesian components.
Figure 2: Shifting of microstates. The position and momentum degrees of freedom of each particle $i$ are displaced according to the map (10) and (11). The vector field ${\boldsymbol{\epsilon}}({\mathbf{r}}_{i})$ parameterizes the transform which is in general spatially inhomogeneous and of a form such that ${\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})$ is a diffeomorphism. The transformation keeps the differential phase space volume element $d{\mathbf{r}}_{i}d{\bf p}_{i}$ conserved for each particle $i$.
We assume that the shifting field ${\boldsymbol{\epsilon}}({\mathbf{r}}_{i})$ has a form that guarantees Eq. (10) to constitute a diffeomorphism. This implies that: i) the transform from ${\mathbf{r}}_{i}$ to ${\mathbf{r}}_{i}^{\prime}$ is smooth, which follows from ${\boldsymbol{\epsilon}}({\mathbf{r}}_{i})$ being smooth, ii) an inverse exists, such that the map from ${\mathbf{r}}_{i}$ to ${\mathbf{r}}_{i}^{\prime}$ is bijective, and iii) the inverse transform from ${\mathbf{r}}_{i}^{\prime}$ to ${\mathbf{r}}_{i}$ is also smooth.
It is straightforward to show that the Jacobian corresponding to the transformation (10) and (11) is unity tschopp2022forceDFT ; hermann2022quantum and hence that the particle-resolved differential phase space volume element $d{\mathbf{r}}_{i}d{\bf p}_{i}$ is preserved under the transformation; see Fig. 2 for an illustration. The variable change expressed by Eqs. (10) and (11) constitutes a canonical transformation in the sense of classical mechanics goldstein2002 . This can be seen by considering a generating function $\cal G$ from which the transformation is obtained by differentiation goldstein2002 . The specific form of this generator is ${\cal G}={\bf p}_{i}^{\prime}\cdot[{\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})]$ tschopp2022forceDFT , where the prime denotes the transformed variables. Using the generic transformation equations ${\mathbf{r}}_{i}^{\prime}=\partial{\cal G}/\partial{\bf p}_{i}^{\prime}$ and ${\bf p}_{i}=\partial{\cal G}/\partial{\mathbf{r}}_{i}$ and solving for the primed variables gives Eqs. (10) and (11).
Alternatively, one can verify that Eqs. (10) and (11) constitute a canonical transformation by explicitly computing the following Poisson bracket identities:
$\{{\mathbf{r}}^{\prime}_{i},{\bf p}^{\prime}_{j}\}$ $=\delta_{ij}{\mathbbm{1}},$ (12)
$\{{\mathbf{r}}^{\prime}_{i},{\mathbf{r}}^{\prime}_{j}\}$ =$=\{{\bf p}^{\prime}_{i},{\bf p}^{\prime}_{j}\}=0,$ (13)
[…]
As a comment on the generality of the phase space shifting given by the joint transformation (10) and (11), this map is constructed to affect individual particles $i$ in an identical way, which is in keeping with the aim of describing the statistical mechanics of identical particles. The position transform (10) is a general smooth map (diffeomorphism) from ${\mathbf{r}}_{i}$ to ${\mathbf{r}}_{i}^{\prime}$. Hence the particular additive form ${\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})$ poses no intrinsic restrictions. Note that any map ${\mathbf{r}}_{i}\to{\mathbf{r}}_{i}^{\prime}({\mathbf{r}}_{i})$ can be rewritten trivially in the form of Eq. (10) as ${\mathbf{r}}_{i}\to{\mathbf{r}}_{i}+[{\mathbf{r}}_{i}^{\prime}({\mathbf{r}}_{i})-{\mathbf{r}}_{i}]$, where the term in brackets then constitutes the shifting field, ${\boldsymbol{\epsilon}}({\mathbf{r}}_{i})={\mathbf{r}}_{i}^{\prime}({\mathbf{r}}_{i})-{\mathbf{r}}_{i}$. Given the choice of position transform (10), the momentum transform (11) then follows uniquely from imposing that the transformation is canonical and hence needs to satisfy Eqs. (12) and (13). These differential equations leave over free remaining integration constants, which are uniquely determined by the additional requirement that the identity transformation is recovered in the case of vanishing shifting field, ${\boldsymbol{\epsilon}}({\mathbf{r}}_{i})=0$.
III.2 Finite phase space shifting
We first demonstrate that chaining two shifting operations constitutes again a single shifting operation. Although this might seem obvious based on mere geometric intuition, carrying out a direct verification based on the structure of transformation equations (10) and (11) is worthwhile. We hence consider a second shifting transform, which is parameterized by a new shifting field ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}^{\prime})$ and which acts on the already transformed variables ${\mathbf{r}}_{i}^{\prime}$ and ${\bf p}_{i}^{\prime}$ that result from applying Eqs. (10) and (11). The result of the consecutive shifting is:
${\mathbf{r}}_{i}^{\prime\prime}$ $={\mathbf{r}}_{i}^{\prime}+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}^{\prime}),$ (14)
${\bf p}_{i}^{\prime\prime}$ $=[{\mathbbm{1}}+\nabla_{i}^{\prime}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}^{\prime})]^{-1}\cdot{\bf p}_{i}^{\prime},$ (15)
where $\nabla_{i}^{\prime}$ denotes the derivative with respect to ${\mathbf{r}}_{i}^{\prime}$. Replacing on the right hand sides of the transformation equations (14) and (15) the variables ${\mathbf{r}}_{i}^{\prime}$ and ${\bf p}_{i}^{\prime}$ via the transformations (10) and (11) yields the following composite expressions:
${\mathbf{r}}_{i}^{\prime\prime}$ $={\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})),$ (16)
${\bf p}_{i}^{\prime\prime}$ $=[{\mathbbm{1}}+\nabla_{i}^{\prime}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))]^{-1}\cdot[{\mathbbm{1}}+\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})]^{-1}\cdot{\bf p}_{i}.$ (17)
We have kept the primed derivative in $\nabla_{i}^{\prime}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))$ in Eq. (17) to indicate the standard gradient of the given vector field with respect to its argument, evaluated at position ${\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$, i.e., $\nabla_{i}^{\prime}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}^{\prime})|_{{\mathbf{r}}_{i}^{\prime}={\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})}$
The conjecture that the mapping (16) and (17) is again a shifting transform implies that a corresponding shifting field ${\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})$ exists, which performs the composite transform in a single step according to:
${\mathbf{r}}_{i}^{\prime\prime}$ $={\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i}),$ (18)
${\bf p}_{i}^{\prime\prime}$ $=[1+\nabla_{i}{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})]^{-1}\cdot{\bf p}_{i}.$ (19)
Comparing the right hand side of the composite position transform (16) with the right hand side of the corresponding single-step position transform (18) allows one to identify:
${\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})$ $={\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})).$ (20)
Equation (20) is an explicit expression for the specific form of the shifting field ${\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})$ that represents the composite transform.
It thus remains to demonstrate the consistency of the single-step momentum transform (19) with the form (20) of the shifting field. Consistency implies that the right hand side of Eqs. (17) and (19) need to be identical, which can only be true in general provided that the following matrices are identical:
$[{\mathbbm{1}}+\nabla_{i}{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})]^{-1}=$
$\qquad[{\mathbbm{1}}+\nabla_{i}^{\prime}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))]^{-1}\cdot[{\mathbbm{1}}+\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})]^{-1}.$ (21)
The validity of Eq. (21) can be verified straightforwardly by inverting the matrices on both its sides and using the chain rule as follows. Keeping in mind to interchange the order of the matrix product upon matrix inversion gives:
${\mathbbm{1}}+\nabla_{i}{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})=[{\mathbbm{1}}+\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})]\cdot[{\mathbbm{1}}+\nabla_{i}^{\prime}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))].$ (22)
Multiplying out the right hand side and simplifying, we can re-write Eq. (22) as $\nabla_{i}{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})=\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+\nabla_{i}^{\prime}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))+[\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})]\cdot\nabla_{i}^{\prime}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))$. This result is identically obtained by building the gradient of Eq. (20) and using the chain rule on the right hand side. This completes our proof that chaining two shifting transformations parameterized by ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}})$ and ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}})$ reduces to a single shifting transformation with shifting field ${\boldsymbol{\epsilon}}_{21}({\mathbf{r}})$ given by Eq. (20).
Figure 3: Group structure of local gauge transformations in statistical mechanics. Rich group structure emerges when chaining multiple local shifting operations together. Appying two consecutive transformations ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ and ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})$ results in a new transformation which is parameterized by ${\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})={\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))$ according to Eq. (20). The final destination point is identical such that ${\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})={\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))$. For given shifting field ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ the corresponding inverse shifting field satisfies ${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))=-{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ according to Eq. (23). The phase space volume element $d{\mathbf{r}}_{i}d{\bf p}_{i}$ is conserved under the transformation, despite its changing shape.
It remains to show the existence of an inverse transform, which is also straightforward. For given ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ the inverse transform ${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})$ is obtained in the concatenation (20) upon requiring that ${\boldsymbol{\epsilon}}_{-11}({\mathbf{r}}_{i})=0$, which constitutes the neutral group element (identity) and we have replaced the index $2\to-1$. Hence from Eq. (20) we obtain the inversion condition
${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))=-{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}),$ (23)
which is an implicit equation for the vector field ${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})$ for given form of ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$. We have to restrict to vector fields ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ that allow ${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})$ to exist, which is guaranteed from the assumption that ${\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})$ constitutes a diffeomorphism, as this implies the existence of an inverse. A simple counterexample is ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})=-{\mathbf{r}}_{i}$ hermann2022quantum which according to Eq. (10) maps all ${\mathbf{r}}_{i}$ to ${\mathbf{r}}_{i}^{\prime}=0$ and is hence trivially not bijective. For completeness, we exchange indices $1$ and $-1$ in Eq. (23) to obtain the equivalent condition ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i}))=-{\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})$. Figure 3 depicts an illustration of both the chaining and the inversion of shifting.
Composing together more than two shifting operations is associative, as is a requirement for valid group structure. An illustration is shown in Fig. 4. Algebraically one has to verify that for three consecutive transforms, as are respectively parameterized by three shifting fields ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$, ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})$, and ${\boldsymbol{\epsilon}}_{3}({\mathbf{r}}_{i})$, the order of the grouping of pairs is irrelevant for the final result. We recall that the non-trivial geometrical structure arises from evaluating a subsequent shift to phase space points that have already been displaced by the prior shift. We demonstrate the equivalence of both groupings explicitly in the following. First, when applying a third transformation to the composite (20), one obtains ${\mathbf{r}}_{i}^{\prime\prime\prime}={\mathbf{r}}_{i}^{\prime\prime}+{\boldsymbol{\epsilon}}_{3}({\mathbf{r}}_{i}^{\prime\prime})={\mathbf{r}}_{i}^{\prime\prime}+{\boldsymbol{\epsilon}}_{3}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i}))={\mathbf{r}}_{i}^{\prime}+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}^{\prime})+{\boldsymbol{\epsilon}}_{3}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})))$, which can be further made explicit as ${\mathbf{r}}_{i}^{\prime\prime\prime}={\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))+{\boldsymbol{\epsilon}}_{3}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})))$. On the other hand, composing the second and third shift gives the displacement vector ${\boldsymbol{\epsilon}}_{32}({\mathbf{r}}_{i}^{\prime})={\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}^{\prime})+{\boldsymbol{\epsilon}}_{3}({\mathbf{r}}_{i}^{\prime}+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}^{\prime}))$ as obtained from Eq. (20) by replacing $1\to 2$, $2\to 3$, and ${\mathbf{r}}_{i}\to{\mathbf{r}}_{i}^{\prime}$. The overall transformed vector is then ${\mathbf{r}}_{i}^{\prime\prime\prime}={\mathbf{r}}_{i}^{\prime}+{\boldsymbol{\epsilon}}_{32}({\mathbf{r}}_{i}^{\prime})={\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{32}({\mathbf{r}}_{i}^{\prime})={\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))+{\boldsymbol{\epsilon}}_{3}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})))$, where we have first replaced ${\mathbf{r}}_{i}^{\prime}$ according to Eq. (10) and have used the prior expression for ${\boldsymbol{\epsilon}}_{32}({\mathbf{r}}_{i}^{\prime})$. Both expressions are identical, which proves associativity for both the position and momentum parts of the phase space shifting. Figure 4 depicts an illustration of the associative shifting structure.
Despite the fact that composition, inversion and associativity are all relatively straightforward, the chaining of two finite shifting operations, as specified by the composite shifting field Eq. (20), is not a commutative operation. This can be seen by interchanging the first and second shifting fields on the right hand side of Eq. (20), which we reproduce for convenience: ${\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})={\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))$. The result is a composite shift given by ${\boldsymbol{\epsilon}}_{12}({\mathbf{r}}_{i})={\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}))$, as obtained from interchanging the indices $1\leftrightarrow 2$. In general the two result will be different from each other, ${\boldsymbol{\epsilon}}_{12}({\mathbf{r}})\neq{\boldsymbol{\epsilon}}_{21}({\mathbf{r}})$; see Fig. 5 for an illustration. We demonstrate in the following Sec. III.3 that the non-commutative character persists for infinitesimal shifting.
In summary, we have shown that locally resolved shifting transformations on phase space given by Eqs. (10) and (11) constitute a non-commutative group. We recall from Sec. II the following requirements: i) the composite of two group element is again a group element, ii) the existence of a neutral element, iii) the existence of an inverse, and iv) associativity. The group elements are parameterized by the form of their corresponding shifting vector field ${\boldsymbol{\epsilon}}({\mathbf{r}})$, which in general can be of finite magnitude.
Figure 4: Illustration of the associativity of the shifting transformation. Shown are different ways of shifting from ${\mathbf{r}}_{i}$ to ${\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{321}({\mathbf{r}}_{i})$. Verifying the associative property implies to show that the combined shifting vector ${\boldsymbol{\epsilon}}_{321}({\mathbf{r}}_{i})$ of three consecutive transformations has two equivalent forms ${\boldsymbol{\epsilon}}_{32}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})={\boldsymbol{\epsilon}}_{3}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i}))+{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})$ (indicated by the two curved arrows). This identity is straightforward to prove from applying Eq. (20) to the two pairs of consecutive operations, which respectively yield ${\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})={\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))$ and ${\boldsymbol{\epsilon}}_{32}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))={\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))+{\boldsymbol{\epsilon}}_{3}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})))$. Inserting these expressions proves the associativity of shifting.
III.3 Infinitesimal position shifting
It is interesting to consider the infinitesimal version of the shifting transformation and in particular to investigate the algebraic structure that emerges from composite shifting. We first address the inversion operation which to first order in the shifting field and its gradient is given by
${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})$ $=-{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}).$ (24)
A derivation of Eq. (24) can be based on expanding the left hand side of the inversion condition (23) to obtain ${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}{\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})=-{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$, which upon re-ordering gives ${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})=-{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot[{\mathbbm{1}}+\nabla_{i}{\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})]$ and ${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})\cdot[{\mathbbm{1}}+\nabla_{i}{\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})]^{-1}=-{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$. Expanding the inverse matrix and truncating after the linear term gives $[{\mathbbm{1}}+\nabla_{i}{\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})]^{-1}={\mathbbm{1}}-\nabla_{i}{\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})$ which yields Eq. (24) upon interchanging ${\boldsymbol{\epsilon}}_{-1}({\mathbf{r}}_{i})$ and ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$.
We consider the action of the spatial displacement (10) on a position-dependent function $f({\mathbf{r}}_{i})$, which is hence moved to $f({\mathbf{r}}_{i}^{\prime})=f({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))$, as obtained from using the position transform (10). Taylor expanding in ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ yields $f({\mathbf{r}}_{i}^{\prime})=f({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}f({\mathbf{r}}_{i})$ to linear order in ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$. From the structure of the Taylor expansion we can identify the differential operator ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}$ as performing the change of the function $f({\mathbf{r}}_{i})$ to $f({\mathbf{r}}_{i}^{\prime})$ to first order in ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$. Correspondingly, shifting by a vector field ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})$ is represented by the operator ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}$.
Successively applying two shifts requires to first expand $f({\mathbf{r}}_{i}^{\prime\prime})=f({\mathbf{r}}_{i}^{\prime}+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}^{\prime}))$ around ${\mathbf{r}}_{i}^{\prime}$ and then, after inserting ${\mathbf{r}}_{i}^{\prime}={\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$, to expand around ${\mathbf{r}}_{i}$. The overall result is $f({\mathbf{r}}_{i}^{\prime\prime})=[f({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}f({\mathbf{r}}_{i})]+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}[f({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}f({\mathbf{r}}_{i})]$, which we re-group as:
$f({\mathbf{r}}_{i}^{\prime\prime})$ $=f({\mathbf{r}}_{i})+[{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})]\cdot\nabla_{i}f({\mathbf{r}}_{i})$
$\quad+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}[{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}f({\mathbf{r}}_{i})].$ (25)
Here the order of the two consecutive shifts is imprinted in the specific form of the last term of the sum on the right hand side of Eq. (25). Reversing the order of the two shifts leads to a corresponding term ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}[{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}f({\mathbf{r}}_{i})]$, which in general is different from the last term in Eq. (25).
Figure 5: Illustration of the non-commutative nature of the shifting transformation. Reversing the order of the two shifting operations that are respectively parameterized by ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ and ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})$ creates a spatial mismatch (dotted arrow) between the two destination points ${\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{12}({\mathbf{r}}_{i})$ and ${\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})$. The non-commutativity is due to the fact that in general ${\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})\neq{\boldsymbol{\epsilon}}_{12}({\mathbf{r}}_{i})$. Here the combined shifting vectors are given by ${\boldsymbol{\epsilon}}_{21}({\mathbf{r}}_{i})={\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}))$ according to Eq. (20) and by ${\boldsymbol{\epsilon}}_{12}({\mathbf{r}}_{i})={\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})+{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i}))$ as obtained by exchanging the indices $1\leftrightarrow 2$.
To systematically capture the emerging difference it is appropriate to consider the commutator of two shifting differential operators ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}$ and ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}$. The following structure is obtained from straightforward calculation:
$[{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i},{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}]$ $={\boldsymbol{\epsilon}}_{\Delta}({\mathbf{r}}_{i})\cdot\nabla_{i},$ (26)
where the brackets denote the commutator of two operators ${\cal O}_{1}$ and ${\cal O}_{2}$ in the usual way as $[{\cal O}_{1},{\cal O}_{2}]={\cal O}_{1}{\cal O}_{2}-{\cal O}_{2}{\cal O}_{1}$. Hence spelling out Eq. (26) explicitly gives $[{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i},{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}]={\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}-{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i}$, with each derivative $\nabla_{i}$ acting on all functions to its right.
The vector field ${\boldsymbol{\epsilon}}_{\Delta}({\mathbf{r}})$ in Eq. (26) that represents the difference between the two composite shifts is given by
${\boldsymbol{\epsilon}}_{\Delta}({\mathbf{r}}_{i})$ $={\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot[\nabla_{i}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})]-{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot[\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})]$ (27)
$=[{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}),{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})]_{L},$ (28)
where the matrices $[\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})]$ and $[\nabla_{i}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})]$ in Eq. (27) are not differential operators but rather tensor fields. In general the right hand side of Eq. (27) will not vanish, ${\boldsymbol{\epsilon}}_{\Delta}({\mathbf{r}}_{i})\neq 0$. The particular combination (27) of two given vector fields ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ and ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})$ constitutes the Lie bracket $[{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}),{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})]_{L}$ as is defined in the standard way robbin2022book in Eq. (27).
Using the Lie bracket (28) we can rewrite the right hand side of the shifting commutator relation (26) to express the commutator relationship as:
$[{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot\nabla_{i},{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot\nabla_{i}]$ $=[{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i}),{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})]_{L}\cdot\nabla_{i},$ (29)
which constitutes a Lie algebra of differential operators.
We have so far formulated the theory on the basis of specific choices of the shifting fields ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}})$, ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}})$ and the resulting form of ${\boldsymbol{\epsilon}}_{\Delta}({\mathbf{r}})$. However, the specific form of these vector fields is not of physical interest, as any observables are independent of the shifting fields, as is consistent with their role as mere gauge functions.
Separating the specificities from the intrinsic nature of the shifting is technically achieved by working with functional differentiation (see Ref. schmidt2022rmp for a practitioners’ account) and we recall that Noether’s theorem addresses such invariant variations noether1918 . Hence having identified the position shifting operators ${\boldsymbol{\epsilon}}({\mathbf{r}}_{i})\cdot\nabla_{i}$, we functionally differentiate with respect to ${\boldsymbol{\epsilon}}({\mathbf{r}})$ as a measure of the response of the shifting operator ${\boldsymbol{\epsilon}}({\mathbf{r}}_{i})\cdot\nabla_{i}$ against changes in the shifting field at a (new) position ${\mathbf{r}}$.
The functional derivative is straightforward to carry out and leads to
$\frac{\delta}{\delta{\boldsymbol{\epsilon}}({\mathbf{r}})}{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})\cdot\nabla_{i}$ $=\delta({\mathbf{r}}_{i}-{\mathbf{r}})\nabla_{i},$ (30)
where we have used that $\delta{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})/\delta{\boldsymbol{\epsilon}}({\mathbf{r}})={\mathbbm{1}}\delta({\mathbf{r}}_{i}-{\mathbf{r}})$ and ${\mathbbm{1}}\cdot\nabla=\nabla$. Multiplying (via a scalar product) the right hand side of Eq. (30) with ${\boldsymbol{\epsilon}}({\mathbf{r}})$ and integrating over ${\mathbf{r}}$ recovers the original position-shifting operator according to:
[…]
Equation (31) can be viewed as a simple case of functional integration (where here the functional integrand is independent of the functional argument ${\boldsymbol{\epsilon}}({\mathbf{r}})$). We refer to Refs. evans1979 ; evans1992 ; sammueller2023whyNeural ; sammueller2024hyperDFT ; sammueller2024whyhyperDFT ; sammueller2024attraction both for descriptions and use of functional integration in more general settings, in particular in the context neural functionals sammueller2023whyNeural ; sammueller2024hyperDFT ; sammueller2024whyhyperDFT ; sammueller2024attraction .
The commutator relationship (29) has a simple bilinear dependence on the two shifting fields. It is hence straightforward to build the mixed second functional derivative of Eq. (26) according to $\delta^{2}/[\delta{\boldsymbol{\epsilon}}_{1}({\mathbf{r}})\delta{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}^{\prime})]$, where ${\mathbf{r}}$ and ${\mathbf{r}}^{\prime}$ are two new position variables. The functional dependence on the shifting fields is explicit on the left hand side of Eq. (29). Functionally differentiating the right hand side of Eq. (29) requires to take into account that ${\boldsymbol{\epsilon}}_{\Delta}({\mathbf{r}}_{i})$ depends on the vector fields ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ and ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})$ via Eq. (27), which gives $\delta^{2}\epsilon_{\Delta c}({\mathbf{r}}_{i})/[\delta\epsilon_{1a}({\mathbf{r}})\delta\epsilon_{2b}({\mathbf{r}}^{\prime})]=\delta({\mathbf{r}}_{i}-{\mathbf{r}})[\nabla_{i}^{a}\delta({\mathbf{r}}_{i}-{\mathbf{r}}^{\prime})\delta_{bc}]-\delta({\mathbf{r}}_{i}-{\mathbf{r}}^{\prime})[\nabla_{i}^{b}\delta({\mathbf{r}}_{i}-{\mathbf{r}})\delta_{ac}]$, where $a,b,c$ denote the Cartesian components and we have used $\delta\epsilon_{a}({\mathbf{r}}_{i})/\delta\epsilon_{b}({\mathbf{r}})=\delta({\mathbf{r}}_{i}-{\mathbf{r}})\delta_{ab}$. The overall result for the mixed functional derivative of the commutator relationship (29) is:
$[\delta({\mathbf{r}}_{i}-{\mathbf{r}})\nabla_{i},\delta({\mathbf{r}}_{i}-{\mathbf{r}}^{\prime})\nabla_{i}]$
$\;=[\nabla_{i}\delta({\mathbf{r}}_{i}-{\mathbf{r}}^{\prime})]\delta({\mathbf{r}}_{i}-{\mathbf{r}})\nabla_{i}-[\nabla_{i}\delta({\mathbf{r}}_{i}-{\mathbf{r}})]\delta({\mathbf{r}}_{i}-{\mathbf{r}}^{\prime})\nabla_{i}^{\sf T}$ (32)
$\;=\delta({\mathbf{r}}_{i}-{\mathbf{r}}^{\prime})\nabla_{i}[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]+[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]\delta({\mathbf{r}}_{i}-{\mathbf{r}})\nabla_{i},$ (33)
where as before $\nabla$ indicates the derivative with respect to ${\mathbf{r}}$ and the superscript $\sf T$ denotes matrix transposition, here of a dyadic product.
We introduce an integral over ${\mathbf{r}}^{\prime\prime}$ on the right hand to express the result as the following operator identity
$[\delta({\mathbf{r}}_{i}-{\mathbf{r}})\nabla_{i},\delta({\mathbf{r}}_{i}-{\mathbf{r}}^{\prime})\nabla_{i}]$
[…]
where ${\sf D}$ is a third-rank tensor that plays the role of Lie structure constants. The $abc$-component of ${\sf D}$ is given by
$D_{abc}({\mathbf{r}},{\mathbf{r}}^{\prime},{\mathbf{r}}^{\prime\prime})$ $=\delta_{bc}\delta({\mathbf{r}}^{\prime\prime}-{\mathbf{r}})[\nabla^{\prime\prime}_{a}\delta({\mathbf{r}}^{\prime\prime}-{\mathbf{r}}^{\prime})]$
$\quad-\delta_{ac}\delta({\mathbf{r}}^{\prime\prime}-{\mathbf{r}}^{\prime})[\nabla^{\prime\prime}_{b}\delta({\mathbf{r}}^{\prime\prime}-{\mathbf{r}})],$ (35)
where $\nabla^{\prime\prime}$ denotes the derivative with respect to ${\mathbf{r}}^{\prime\prime}$.
In summary we have demonstrated for the case of position-only shifting the relationship of the spatial displacement with the action of shifting differential operators. We have shown that the non-commutative nature that is inherent when considering finite displacements persists when Taylor expanding and keeping only the lowest nontrivial terms. The non-commutative aspect is then reflected by algebraic relationships. Specifically, when expressed in the form (35) it is apparent that the particular form of the product of the delta distribution and derivative operator, $\delta({\mathbf{r}}_{i}-{\mathbf{r}})\nabla_{i}$, persists as a result of building the commutator.
In Sec. IV.3, we demonstrate that much of this favourable structure can be retained when considering the joint phase space transformation of coordinates and momenta. We recall that this constitutes a canonical transformation, which the present position-only shifting with unchanged momenta does not.
IV Statistical Mechanics
IV.1 Partition sum and thermal averages
[…]
$\langle\cdot\rangle={\rm Tr}\,\cdot{\rm e}^{-\beta H({\mathbf{r}}^{N},{\bf p}^{N})}/Z.$ (36)
More explicitly the thermal average of an observable $\hat{A}({\mathbf{r}}^{N})$ is given by
[…]
We first rewrite the phase space integral in Eq. (37) in an identical way upon merely renaming the unprimed by primed variables:
[…]
where the primed partition sum $Z^{\prime}$ is expressed in the primed variables and $Z^{\prime}=Z$.
We next perform the canonical variable transformation given by Eqs. (10) and (11). The phase space volume element is thereby conserved such that $d{\mathbf{r}}_{i}d{\bf p}_{i}=d{\mathbf{r}}_{i}^{\prime}d{\bf p}_{i}^{\prime}$ for each particle $i$. The result is
[…]
$=A[{\boldsymbol{\epsilon}}],$ (40)
The dependence of the integrand in Eq. (39) on the unprimed variables is suppressed in the notation for clarity and it occurs via the transformation equations (10) and (11) that render the ${\mathbf{r}}_{i}^{\prime}$ and ${\bf p}_{i}^{\prime}$ parametrically dependent on ${\mathbf{r}}_{i}$ and ${\bf p}_{i}$ and functionally dependent on the form of the shifting field ${\boldsymbol{\epsilon}}({\mathbf{r}})$. Spelling out these dependencies explicitly we have ${\mathbf{r}}_{i}^{\prime}={\mathbf{r}}_{i}^{\prime}({\mathbf{r}}_{i},[{\boldsymbol{\epsilon}}])$ according to Eq. (10) and ${\bf p}_{i}^{\prime}={\bf p}^{\prime}({\mathbf{r}}_{i},{\bf p}_{i},[{\boldsymbol{\epsilon}}])$ according to Eq. (11).
After the unprimed phase space integral is performed, at face value Eq. (39) depends functionally on the shifting field, which we have indicated as $A[{\boldsymbol{\epsilon}}]$ in Eq. (40). Comparing Eq. (37) with Eq. (40) gives the invariance condition $A=A[{\boldsymbol{\epsilon}}]$, where the left hand side can also be understood as $A[{\boldsymbol{\epsilon}}]$ being evaluated at vanishing shifting field, ${\boldsymbol{\epsilon}}({\mathbf{r}})=0$, such that $A[0]=A[{\boldsymbol{\epsilon}}]$.
As the phase space shifting affects systems with fixed number of particles, the argumentation carries over to the grand ensemble where each system with $N$ particles is displaced identically. Hence besides the changes in the definition of the grand ensemble averages, no structural repercussions occur for any of the results.
We consider Hamiltonians that consist of kinetic, interparticle, and external energy contributions according to the sum:
$H({\mathbf{r}}^{N},{\bf p}^{N})$ $=\sum_{i}\frac{{\bf p}_{i}^{2}}{2m}+u({\mathbf{r}}^{N})+\sum_{i}V_{\rm ext}({\mathbf{r}}_{i}),$ (41)
where the sums run over all $N$ particle indices $i$, $m$ denotes the particle mass, $u({\mathbf{r}}^{N})$ is the interparticle interaction potential, and $V_{\rm ext}({\mathbf{r}})$ is an external one-body potential.
IV.2 One-body observables
We lay out the relevant one-body observables, which are pertinent in the equilibrium sum rules that follow from the gauge invariance as described above. The one-body level of correlation functions is integral in the classical density functional formulation of statistical mechanics evans1979 ; evans1992 ; evans2016 ; hansen2013 ; schmidt2022rmp and it naturally extends to the dynamical problems, see Ref. schmidt2022rmp for systematic derivations on the basis of classical Hamiltonian, overdamped Brownian, and quantum many-body time evolution.
The one-body localization enables one to carry out systematic coarse-graining in the sense of reduction of the full phase space information, while it remains microscopically sharp, such that all features on microscopic length scales are resolved without any principal loss. The primary one-body observable is the one-body density distribution or density “profile” $\rho({\mathbf{r}})$. When working with functional relationships two- and higher-body observables are in principle accessible via functional differentiation both in evans1979 ; evans1992 ; evans2016 ; hansen2013 ; schmidt2022rmp and out-of-equilibrium schmidt2022rmp .
As an aside, to reach beyond the hierarchy of increasingly higher-order correlation functions, the recent hyperdensity functional theory sammueller2024hyperDFT ; sammueller2024whyhyperDFT allows one to treat complex order parameters, including algorithmically defined phase space functions, on the basis of the density profile as the fundamental variable, as is ensured by the Mermin-Evans mermin1965 ; evans1979 density functional map.
The response of the Hamiltonian $H$ against changes in the form of the external potential $V_{\rm ext}({\mathbf{r}})$ is measured by the functional derivative $\hat{\rho}({\mathbf{r}})=\delta H/\delta V_{\rm ext}({\mathbf{r}})$, where the density “operator” (phase space function) is obtained as:
$\hat{\rho}({\mathbf{r}})$ $=\sum_{i}\delta({\mathbf{r}}-{\mathbf{r}}_{i}).$ (42)
The density profile is the thermal average, $\rho({\mathbf{r}})=\langle\hat{\rho}({\mathbf{r}})\rangle$. Its use allows one to express the average localized external force density as $-\rho({\mathbf{r}})\nabla V_{\rm ext}({\mathbf{r}})$, where we recall that $V_{\rm ext}({\mathbf{r}})$ is the external potential that a particle situated at position ${\mathbf{r}}$ contributes, and $-\nabla V_{\rm ext}({\mathbf{r}})$ is the corresponding external force field. It is straightforward to show the equality $-\rho({\mathbf{r}})\nabla V_{\rm ext}({\mathbf{r}})=\langle\sum_{i}\delta({\mathbf{r}}-{\mathbf{r}}_{i}){\bf f}_{\rm ext}({\mathbf{r}}_{i})\rangle$, where the external force field ${\bf f}_{\rm ext}({\mathbf{r}})=-\nabla V_{\rm ext}({\mathbf{r}})$.
The localized interparticle force density “operator” (phase space function) is defined as:
$\hat{\bf F}_{\rm int}({\mathbf{r}})$ $=-\sum_{i}\delta({\mathbf{r}}-{\mathbf{r}}_{i})\nabla_{i}u({\mathbf{r}}^{N}).$ (43)
The corresponding mean interparticle force density is simply the average ${\bf F}_{\rm int}({\mathbf{r}})=\langle\hat{\bf F}_{\rm int}({\mathbf{r}})\rangle$, with $\hat{\bf F}_{\rm int}({\mathbf{r}})$ defined by Eq. (43). Interparticle and external force densities can be combined into a dedicated potential force density:
${\bf F}_{U}({\mathbf{r}})$ $={\bf F}_{\rm int}({\mathbf{r}})-\rho({\mathbf{r}})\nabla V_{\rm ext}({\mathbf{r}}),$ (44)
where the subscript $U$ refers to total potential energy, as given as a phase space function by $\hat{U}=u({\mathbf{r}}^{N})+\sum_{i}V_{\rm ext}({\mathbf{r}}_{i})$. The corresponding potential force density “operator” is analogously defined via
$\hat{\bf F}_{U}({\mathbf{r}})$ $=\hat{\bf F}_{\rm int}({\mathbf{r}})-\hat{\rho}({\mathbf{r}})\nabla V_{\rm ext}({\mathbf{r}}),$ (45)
and by construction the average is the mean potential force density, ${\bf F}_{U}({\mathbf{r}})=\langle\hat{\bf F}_{U}({\mathbf{r}})\rangle$.
In equilibrium the potential forces are balanced by a thermal force density which arises from the kinetic energy contribution to the Hamiltonian (41). The general phase space function that underlies this effect is the localized kinetic stress “operator” $\hat{\boldsymbol{\tau}}({\mathbf{r}})$ which is defined as:
$\hat{\boldsymbol{\tau}}({\mathbf{r}})$ $=-\sum_{i}\frac{{\bf p}_{i}{\bf p}_{i}}{m}\delta({\mathbf{r}}-{\mathbf{r}}_{i}),$ (46)
where ${\bf p}_{i}{\bf p}_{i}$ denotes the dyadic product of the momentum of particle $i$ with itself. It is straightforward to show upon using the properties of the Maxwellian, i.e., the Gaussian distribution of the momenta according to the Boltzmann factor, that $\nabla\cdot{\boldsymbol{\tau}}({\mathbf{r}})=\nabla\cdot\langle\hat{\boldsymbol{\tau}}({\mathbf{r}})\rangle=-k_{B}T\nabla\rho({\mathbf{r}})$, which acts as a thermal, diffusive force density.
Collecting kinetic, interparticle, and external contributions, we can define a total force density operator:
$\hat{\bf F}({\mathbf{r}})$ $=\nabla\cdot\hat{\boldsymbol{\tau}}({\mathbf{r}})+\hat{\bf F}_{\rm int}({\mathbf{r}})-\hat{\rho}({\mathbf{r}})\nabla V_{\rm ext}({\mathbf{r}}),$ (47)
where we refer the Reader to Ref. schmidt2022rmp for the derivation from the second time derivative of the density operator (42). In thermal equilibrium the average one-body force density ${\bf F}({\mathbf{r}})=\langle\hat{\bf F}({\mathbf{r}})\rangle$ vanishes,
${\bf F}({\mathbf{r}})$ $=0,$ (48)
which upon spelling out the three individual terms is analogous to the more detailed form:
$-k_{B}T\nabla\rho({\mathbf{r}})+{\bf F}_{\rm int}({\mathbf{r}})-\rho({\mathbf{r}})\nabla V_{\rm ext}({\mathbf{r}})$ $=0.$ (49)
The equilibrium force density balance (48) can be verified by explicit calculation of the thermal averages upon using partial integration on phase space to manipulate the occurring coordinate derivatives. In particular when the interparticle force density ${\bf F}_{\rm int}({\mathbf{r}})$ is expressed as an integral over the two-body density multiplied by the pair force, Eq. (48) is commonly referred to as the Yvon-Born-Green (YBG) equation yvon1935 ; born1946 ; hansen2013 .
Finally, it is useful to define a configurational force density “operator” as
$\beta\hat{\bf F}_{c}({\mathbf{r}})$ $=-\nabla\hat{\rho}({\mathbf{r}})+\beta\hat{\bf F}_{U}({\mathbf{r}}),$ (50)
which on average $\langle\hat{\bf F}_{c}({\mathbf{r}})\rangle=0$, which is identical to the equilibrium force density balance (48). Note that for purely configurational observables $\hat{A}({\mathbf{r}}^{N})$, due to the simplicity of the Maxwellian, we have $\langle\hat{A}({\mathbf{r}}^{N})\hat{\bf F}_{c}({\mathbf{r}})\rangle=\langle\hat{A}({\mathbf{r}}^{N})\hat{\bf F}({\mathbf{r}})\rangle$, see e.g. Ref. hermann2023whatIsLiquid for further details of the thermal average of the kinetic stress operator (46).
IV.3 Sum rules from configuration shifting
The mathematical structure of infinitesimal position shifting described in Sec. III.3 can be directly used to derive statistical mechanical sum rules, as we demonstrate in the following. We define the following localized position-shifting operators
$\boldsymbol{\sigma}_{c}({\mathbf{r}})=\sum_{i}\delta({\mathbf{r}}-{\mathbf{r}}_{i})\nabla_{i},$ (51)
where the subscript $c$ indicates that only the configuration ${\mathbf{r}}^{N}$ is affected, independent of the momentum degrees of freedom ${\bf p}^{N}$.
The adjoint operator is given by
$\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}})$ $=-\boldsymbol{\sigma}_{c}({\mathbf{r}})+\nabla\hat{\rho}({\mathbf{r}}),$ (52)
[…]
The configurational shifting operators at two generic positions ${\mathbf{r}}$ and ${\mathbf{r}}^{\prime}$ satisfy the commutator relationship
$[\boldsymbol{\sigma}_{c}({\mathbf{r}}),\boldsymbol{\sigma}_{c}({\mathbf{r}}^{\prime})]$ $=\boldsymbol{\sigma}_{c}({\mathbf{r}}^{\prime})[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]+[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]\boldsymbol{\sigma}_{c}({\mathbf{r}}),$ (53)
Equation (53) follows straightforwardly from the corresponding single-particle commutator (33) upon summing the latter over $i$, then identifying the explicit form (51) of $\boldsymbol{\sigma}_{c}({\mathbf{r}})$, and noticing that the commutator vanishes for two different particles $i\neq j$. We use the prime in the two different roles to denote the generic position variable ${\mathbf{r}}^{\prime}$, which is in general unrelated to ${\mathbf{r}}$, as well as to indicate the transformed phase space coordinate ${\mathbf{r}}_{i}^{\prime}$, which is related to ${\mathbf{r}}_{i}$ via the position transform (10).
It is interesting to apply the localized configuration shift operator (51) to the Hamiltonian, which yields
$-[\boldsymbol{\sigma}_{c}({\mathbf{r}})H]$ $=\hat{\bf F}_{U}({\mathbf{r}}),$ (54)
where we recall the potential force density “operator” $\hat{\bf F}_{U}({\mathbf{r}})$ as the sum (45) of interparticle and external contributions. The brackets on the left hand side of Eq. (54) indicate the range of action of $\boldsymbol{\sigma}_{c}({\mathbf{r}})$, i.e., it only operates on $H$ but not beyond. The explicit verification of Eq. (54) can be based on the definition (51) of $\boldsymbol{\sigma}_{c}({\mathbf{r}})$, the form (41) of the Hamiltonian, and the definition (44) of the potential force operator $\hat{\bf F}_{U}({\mathbf{r}})$
Similarly, when applied to the Boltzmann factor, the chain rule together with Eq. (54) yields straightforwardly
$[\boldsymbol{\sigma}_{c}({\mathbf{r}}){\rm e}^{-\beta H}]$ $=\beta\hat{\bf F}_{U}({\mathbf{r}}){\rm e}^{-\beta H}.$ (55)
Equation (55) constitutes a fundamental link between spatial shifting of the (equilibrium) many-body probability distribution and the emergence of the localized force density operator. It remains to exploit the specific operator structure of the localized shifting operators, as embodied in Eqs. (52) and (53), and to build the average over the equilibrium ensemble.
To derive hyperforce sum rules we multiply the adjoint identity Eq. (52) from the left by a purely coordinate-dependent phase space function $\hat{A}({\mathbf{r}}^{N})$ and then build the thermal average of the result, which gives:
$\langle\hat{A}\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}})\rangle$ $=-\langle\hat{A}\boldsymbol{\sigma}_{c}({\mathbf{r}})\rangle+\nabla\langle\hat{A}\hat{\rho}({\mathbf{r}})\rangle.$ (56)
Recalling the thermal average being defined via Eq. (36) the left hand side of Eq. (56) is explicitly given by $\langle\hat{A}\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}})\rangle={\rm Tr}\,\hat{A}\boldsymbol{\sigma}^{\dagger}({\mathbf{r}}){\rm e}^{-\beta H}/Z={\rm Tr}\,[\boldsymbol{\sigma}({\mathbf{r}})\hat{A}]{\rm e}^{-\beta H}/Z=\langle[\boldsymbol{\sigma}_{c}\hat{A}]\rangle$, where in the second step we have made use of the defining property of the adjoint operator. Making also the first term on the right hand side of Eq. (56) more explicit yields $-\langle\hat{A}\boldsymbol{\sigma}_{c}({\mathbf{r}})\rangle=-{\rm Tr}\,\hat{A}\boldsymbol{\sigma}_{c}({\mathbf{r}}){\rm e}^{-\beta H}/Z=-{\rm Tr}\,\hat{A}\beta\hat{\bf F}_{U}({\mathbf{r}}){\rm e}^{-\beta H}/Z=-\langle\hat{A}\beta\hat{\bf F}_{U}({\mathbf{r}})\rangle$, where we have first used the chain rule and then applied the configuration shift operator to the Boltzmann factor according to Eq. (55).
As the second term on the right hand side of Eq. (56) is already in the form of an explicit average, we can collect all terms to formulate the following exact equilibrium sum rule
${\bf S}_{A}({\mathbf{r}})$ $=-\langle\hat{A}\beta\hat{\bf F}_{U}({\mathbf{r}})\rangle+\nabla\langle\hat{A}\hat{\rho}({\mathbf{r}})\rangle,$ (57)
where the hyperforce density ${\bf S}_{A}({\mathbf{r}})=\langle\hat{\bf S}_{A}({\mathbf{r}})\rangle$ is given in the present case of mere dependence on coordinates (independet of momenta) as
$\hat{S}_{A}({\mathbf{r}})$ $=[\boldsymbol{\sigma}_{c}({\mathbf{r}})\hat{A}({\mathbf{r}}^{N})]$ (58)
$=\sum_{i}\delta({\mathbf{r}}-{\mathbf{r}}_{i})[\nabla_{i}\hat{A}({\mathbf{r}}^{N})].$ (59)
The form (59) follows directly from the definition of the position shift operator (51). Using the configurational force density operator $\hat{\bf F}_{c}({\mathbf{r}})$ given by Eq. (50) and rearranging allows one to rewrite the sum rule Eq. (57) as
${\bf S}_{A}({\mathbf{r}})+\langle\hat{A}({\mathbf{r}}^{N})\beta\hat{\bf F}_{c}({\mathbf{r}})\rangle$ $=0.$ (60)
The situation becomes even richer when using two (or more) shifting operators due to a larger variety of re-writing of equivalent expressions. This multitude could potentially be useful in concrete applications. We here restrict ourselves to two specific cases and hence first consider the operator algebra (53). To arrive at a sum rule, we multiply Eq. (53) by $\hat{A}({\mathbf{r}}^{N})$ from the left and then build the thermal average. The first term of the commutator is $\langle\hat{A}\boldsymbol{\sigma}_{c}({\mathbf{r}})\boldsymbol{\sigma}_{c}({\mathbf{r}}^{\prime})\rangle=\langle[\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}})\hat{A}]\boldsymbol{\sigma}_{c}({\mathbf{r}}^{\prime})\rangle=\langle[-\boldsymbol{\sigma}_{c}({\mathbf{r}})\hat{A}+\nabla\hat{\rho}({\mathbf{r}})\hat{A}]\beta\hat{\bf F}_{U}({\mathbf{r}}^{\prime})\rangle=-\langle\hat{\bf S}_{A}({\mathbf{r}})\beta\hat{\bf F}_{U}({\mathbf{r}}^{\prime})\rangle+\nabla\langle\hat{A}\hat{\rho}({\mathbf{r}})\beta\hat{\bf F}_{U}({\mathbf{r}}^{\prime})\rangle$, where we have first re-written via the adjoint $\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}})$, used its relationship (52), and the definition (58) of the configurational hyperforce density operator ${\bf S}_{A}({\mathbf{r}})$.
The result for the second term of the commutator follows from an analogous chain of steps or via simply exchanging ${\mathbf{r}}$ and ${\mathbf{r}}^{\prime}$ and transposing, which yields $\langle\hat{A}\boldsymbol{\sigma}_{c}({\mathbf{r}}^{\prime})\boldsymbol{\sigma}_{c}({\mathbf{r}})\rangle^{\sf T}=-\langle\hat{\bf S}_{A}({\mathbf{r}}^{\prime})\beta\hat{\bf F}_{U}({\mathbf{r}})\rangle^{\sf T}+\nabla^{\prime}\langle\hat{A}\hat{\rho}({\mathbf{r}}^{\prime})\beta\hat{\bf F}_{U}({\mathbf{r}})\rangle^{\sf T}$. Averaging the right hand side of the commutator relationship (53) is straightforward and requires the averages $\langle\hat{A}\boldsymbol{\sigma}_{c}({\mathbf{r}}^{\prime})\rangle=\langle\hat{A}\beta\hat{\bf F}_{U}({\mathbf{r}}^{\prime})\rangle$ and $\langle\hat{A}\boldsymbol{\sigma}_{c}({\mathbf{r}})\rangle=\langle\hat{A}\beta\hat{\bf F}_{U}({\mathbf{r}})\rangle$.
Collecting all terms we obtain as a result the following equilibrium hyperforce sum rule:
$-\langle\hat{\bf S}_{A}({\mathbf{r}})\beta\hat{\bf F}_{U}({\mathbf{r}}^{\prime})\rangle+\nabla\langle\hat{A}\rho({\mathbf{r}})\beta\hat{\bf F}_{U}({\mathbf{r}}^{\prime})\rangle$
$+\langle\hat{\bf S}_{A}({\mathbf{r}}^{\prime})\beta\hat{\bf F}_{U}({\mathbf{r}})\rangle^{\sf T}-\nabla^{\prime}\langle\hat{A}\rho({\mathbf{r}}^{\prime})\beta\hat{\bf F}_{U}({\mathbf{r}})\rangle^{\sf T}$
$\quad=\langle\hat{A}\beta\hat{\bf F}_{U}({\mathbf{r}}^{\prime})\rangle[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]+[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]\langle\hat{A}\beta\hat{\bf F}_{U}({\mathbf{r}})\rangle.$ (61)
The right hand side of Eq. (61) can upon using Eq. (60) alternatively be expressed as $[\langle\hat{A}\nabla^{\prime}\hat{\rho}({\mathbf{r}}^{\prime})\rangle-{\bf S}_{A}({\mathbf{r}}^{\prime})][\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]+[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})][\langle\hat{A}\nabla\hat{\rho}({\mathbf{r}})\rangle-{\bf S}_{A}({\mathbf{r}})]$.
As an alternative starting point we use the adjoint of the commutator relation (53), which is given by
$-[\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}}),\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}}^{\prime})]$ $=\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}}^{\prime})[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]$
$\quad+[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}}),$ (62)
Multiplying Eq. (62) by the configuration-dependent observable $\hat{A}({\mathbf{r}}^{N})$ from the left and building the thermal average on both sides involves the following three terms: $-\langle\hat{A}\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}})\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}}^{\prime})\rangle=\langle[\boldsymbol{\sigma}_{c}({\mathbf{r}})\hat{A}][\boldsymbol{\sigma}_{c}({\mathbf{r}}^{\prime})-\nabla^{\prime}\hat{\rho}({\mathbf{r}}^{\prime})]\rangle=\langle\hat{\bf S}_{A}({\mathbf{r}})[\beta\hat{\bf F}_{U}({\mathbf{r}}^{\prime})-\nabla^{\prime}\hat{\rho}({\mathbf{r}}^{\prime})]\rangle=\langle\hat{\bf S}_{A}({\mathbf{r}})\beta\hat{\bf F}_{c}({\mathbf{r}}^{\prime})\rangle$, and similarly $-\langle\hat{A}\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}}^{\prime})\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}})\rangle^{\sf T}=\langle\beta\hat{\bf F}_{c}({\mathbf{r}})\hat{\bf S}_{A}({\mathbf{r}}^{\prime})\rangle$, as well as $\langle\hat{A}\boldsymbol{\sigma}_{c}^{\dagger}({\mathbf{r}})\rangle=\langle[\boldsymbol{\sigma}_{c}({\mathbf{r}})\hat{A}]\rangle=\langle\hat{\bf S}_{A}({\mathbf{r}})\rangle={\bf S}_{A}({\mathbf{r}})$. Collecting these results leads to the following sum rule:
$\langle\hat{\bf S}_{A}({\mathbf{r}})\beta\hat{\bf F}_{c}({\mathbf{r}}^{\prime})\rangle-\langle\beta\hat{\bf F}_{c}({\mathbf{r}})\hat{\bf S}_{A}({\mathbf{r}}^{\prime})\rangle$
$={\bf S}_{A}({\mathbf{r}}^{\prime})[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]+[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]{\bf S}_{A}({\mathbf{r}}),$ (63)
which is the analogous position-dependent version of the more general (momentum-dependent) result of Ref. mueller2024gauge , which features the momentum-dependent force operator $\hat{\bf F}({\mathbf{r}})$, as given by (47), in lieu of its configurational counterpart $\hat{\bf F}_{c}({\mathbf{r}})$, as given by Eq. (50).
The examples (57), (61) and (63) demonstrate that an algebraic structure of correlation functions emerges from both the adjoint property (52) and the commutator (53). We have restricted ourselves to considering products of two shifting operators. The present shifting in particle positions does not yet constitute a canonical transform and we next demonstrate that incorporating the momentum transform allows one to obtain canonical shifting transformations which retain the useful commutator structure revealed above for purely configurational shifting.
Figure 6: Demonstration of particle gauge invariance via Monte Carlo simulations. Hard rod particles of size $a$ confined between two hard walls, as represented by a scaled external potential $\beta V_{\rm ext}({\mathbf{r}})$ (gray lines) that vanishes for $0.5a\leq x\leq 9.5a$ and is $\infty$ otherwise. Results are shown for the unshifted system (left column), a shifted system (middle column), and a corrupted system (right column) with amplitude of the shifting field beyond the threshold of invertibility, such that the transformation fails to be bijective and is hence no longer a gauge transformation. Shown is the scaled sinusoidal shifting field $\epsilon(x)/a$ as a function of scaled position $x/a$ (top panels), the scaled density profile $\rho(x)a$ (middle panels), and the scaled one-body phase space distribution $f(x,p)a\sqrt{mk_{B}T}$ (bottom panels), where the variables are the scaled position $x/a$ and scaled momentum $p/\sqrt{mk_{B}T}$. The shifting transformation displaces both coordinates and momenta. Despite the difference in Monte Carlo simulation moves and acceptance events, the results in the original system (left column) and in the transformed sytem (middle column) are numerically identical to each other, while the invalidly transformed system develops artifacts (right column).
IV.4 Infinitesimal phase space shifts and Lie algebra
We are now in a position to describe the consequence of the full phase space shifting transformation mueller2024gauge , as represented on the level of microstates by the maps that displace particle positions and momentum degrees of freedom according to Eqs. (10) and (11), respectively. Corresponding localized shifting operators displace in position in analogy to the configurational shifting operators $\boldsymbol{\sigma}_{c}({\mathbf{r}})$, as defined in Eq. (51).
Specifically, the full position-resolved phase space shifting operators are defined as mueller2024gauge :
$\boldsymbol{\sigma}({\mathbf{r}})$ $=\sum_{i}\big{[}\delta({\mathbf{r}}-{\mathbf{r}}_{i})\nabla_{i}+{\bf p}_{i}\nabla\delta({\mathbf{r}}-{\mathbf{r}}_{i})\cdot\nabla_{{\bf p}_{i}}\big{]}.$ (64)
When applied to a purely configuration-dependent observable $\hat{A}({\mathbf{r}}^{N})$, the dependence on momentum vanishes and $\boldsymbol{\sigma}({\mathbf{r}})\hat{A}({\mathbf{r}}^{N})=\boldsymbol{\sigma}_{c}({\mathbf{r}})\hat{A}({\mathbf{r}}^{N})$, where $\boldsymbol{\sigma}_{c}({\mathbf{r}})$ is defined in Eq. (51). The momentum contribution to the transform [second term in the sum on the right hand side of Eq. (64)] constitutes spatially localized momentum shifting operators with the explicit form:
$\boldsymbol{\sigma}_{p}({\mathbf{r}})$ $=\sum_{i}{\bf p}_{i}\nabla\delta({\mathbf{r}}-{\mathbf{r}}_{i})\cdot\nabla_{{\bf p}_{i}}.$ (65)
Together with the configurational shifting operator $\boldsymbol{\sigma}_{c}({\mathbf{r}})$ we can express $\boldsymbol{\sigma}({\mathbf{r}})$ as given via Eq. (64) as the sum of configurational and momentum contributions according to:
$\boldsymbol{\sigma}({\mathbf{r}})$ $=\boldsymbol{\sigma}_{c}({\mathbf{r}})+\boldsymbol{\sigma}_{p}({\mathbf{r}}).$ (66)
The adjoint momentum shift operator is given as:
$\boldsymbol{\sigma}_{p}^{\dagger}({\mathbf{r}})$ $=-\boldsymbol{\sigma}_{p}({\mathbf{r}})-\nabla\hat{\rho}({\mathbf{r}}),$ (67)
which follows from integration by parts on momentum space, using $\nabla\cdot\nabla_{{\bf p}_{i}}{\bf p}_{i}=\nabla\cdot{\mathbbm{1}}=\nabla$, and identifying the density operator $\hat{\rho}({\mathbf{r}})$, see its definition (42).
The adjoint of the full phase space shifting operator $\boldsymbol{\sigma}({\mathbf{r}})$ is then the sum of the individual adjoint identities (52) and (67). The result is the simple anti-self adjoint property:
$\boldsymbol{\sigma}^{\dagger}({\mathbf{r}})=-\boldsymbol{\sigma}({\mathbf{r}}).$ (68)
Despite an increase in apparent structural complexity of $\boldsymbol{\sigma}({\mathbf{r}})$ over $\boldsymbol{\sigma}_{c}({\mathbf{r}})$, the corresponding commutator relationship remains simple:
$[\boldsymbol{\sigma}({\mathbf{r}}),\boldsymbol{\sigma}({\mathbf{r}}^{\prime})]$ $=\boldsymbol{\sigma}({\mathbf{r}}^{\prime})[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]+[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]\boldsymbol{\sigma}({\mathbf{r}}),$ (69)
which is identical in form to the corresponding configurational identity (53).
The following two general properties are straightforward to show:
$[\boldsymbol{\sigma}({\mathbf{r}}),\boldsymbol{\sigma}({\mathbf{r}}^{\prime})]^{\dagger}$ $=-[\boldsymbol{\sigma}({\mathbf{r}}),\boldsymbol{\sigma}({\mathbf{r}}^{\prime})],$ (70)
$[\boldsymbol{\sigma}({\mathbf{r}}),\boldsymbol{\sigma}({\mathbf{r}}^{\prime})]$ $=-[\boldsymbol{\sigma}({\mathbf{r}}^{\prime}),\boldsymbol{\sigma}({\mathbf{r}})]^{\sf T},$ (71)
where Eq. (70) holds for two self-adjoint operators and Eq. (71) is valid for two vectorial operators. Furthermore the Jacobi identity holds:
$[\sigma_{a}({\mathbf{r}}),[\sigma_{b}({\mathbf{r}}^{\prime}),\sigma_{c}({\mathbf{r}}^{\prime\prime})]]\quad$
$+[\sigma_{b}({\mathbf{r}}^{\prime}),[\sigma_{c}({\mathbf{r}}^{\prime\prime}),\sigma_{a}({\mathbf{r}})]]$
$+[\sigma_{c}({\mathbf{r}}^{\prime\prime}),[\sigma_{a}({\mathbf{r}}),\sigma_{b}({\mathbf{r}}^{\prime})]]$ $=0.$ (72)
Applying the localized phase space shifting operators to the Hamiltonian and to the Boltzmann factor respectively yields:
$-[\boldsymbol{\sigma}({\mathbf{r}})H]$ $=\hat{\bf F}({\mathbf{r}}),$ (73)
$[\boldsymbol{\sigma}({\mathbf{r}}){\rm e}^{-\beta H}]$ $=\beta\hat{\bf F}({\mathbf{r}}){\rm e}^{-\beta H},$ (74)
where the total force density operator, including the divergences of the explicit momentum-dependent kinetic stress tensor $\nabla\cdot\hat{\boldsymbol{\tau}}({\mathbf{r}})$, is given by Eq. (47). Equations (73) and (74) mirror the corresponding relationships (54) and (55) for configurational shifting on the basis of $\boldsymbol{\sigma}_{c}({\mathbf{r}})$ given by Eq. (51) and the potential force density $\hat{\bf F}_{U}({\mathbf{r}})$ given by Eq. (45).
The procedure of obtaining sum rules from full phase space shifting is similar in structure to working with configurational shifts described in Sec. IV.3. Application to a phase space function $\hat{A}({\mathbf{r}}^{N},{\bf p}^{N})$ with full configurational dependence on ${\mathbf{r}}^{N}$ and momentum dependence on ${\bf p}^{N}$ yields robitschko2024any ; mueller2024gauge :
$\hat{\bf S}_{A}({\mathbf{r}})$ $=[\boldsymbol{\sigma}({\mathbf{r}})\hat{A}]$ (75)
$=\sum_{i}\delta({\mathbf{r}}-{\mathbf{r}}_{i})(\nabla_{i}\hat{A})$
$\quad+\nabla\cdot\sum_{i}\delta({\mathbf{r}}-{\mathbf{r}}_{i})(\nabla_{{\bf p}_{i}}\hat{A}){\bf p}_{i}.$ (76)
Application in the form $\langle[\boldsymbol{\sigma}^{\dagger}({\mathbf{r}})\hat{A}]\rangle=\langle\hat{A}\boldsymbol{\sigma}({\mathbf{r}})\rangle$ yields upon using the self-adjoint property (68) and the generation of the force density operator via application to the Boltzmann factor (74) the one-body hyperforce sum rule:
${\bf S}_{A}({\mathbf{r}})+\langle\hat{A}\beta\hat{\bf F}({\mathbf{r}})\rangle$ $=0,$ (77)
which applies to general observables $\hat{A}({\mathbf{r}}^{N},{\bf p}^{N})$, with the one-body hyperforce density being the average ${\bf S}_{A}({\mathbf{r}})=\langle\hat{\bf S}_{A}({\mathbf{r}})\rangle$ of the corresponding hyperforce phase space function (76), and the full force density operator $\hat{\bf F}({\mathbf{r}})$ given by Eq. (47). Again we point out the formal similarity with the configuration versions (57) and (60).
Averaging the commutator relationship (69) yields the following hyperforce sum rule:
$\langle\hat{\bf S}_{A}({\mathbf{r}})\beta\hat{\bf F}({\mathbf{r}}^{\prime})\rangle-\langle\beta\hat{\bf F}({\mathbf{r}})\hat{\bf S}_{A}({\mathbf{r}}^{\prime})\rangle$
$\qquad={\bf S}_{A}({\mathbf{r}}^{\prime})[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]+[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]{\bf S}_{A}({\mathbf{r}}),$ (78)
which is in analogy to the configurational version (63). The right hand side of Eq. (78) vanishes for ${\mathbf{r}}\neq{\mathbf{r}}^{\prime}$ and the following exchange symmetry holds for the case of distinct positions:
$\langle\hat{\bf S}_{A}({\mathbf{r}})\hat{\bf F}({\mathbf{r}}^{\prime})\rangle$ $=\langle\hat{\bf F}({\mathbf{r}})\hat{\bf S}_{A}({\mathbf{r}}^{\prime})\rangle.$ (79)
It remains to formulate the Lie algebra for phase space shifting. For prescribed form of the shifting field ${\boldsymbol{\epsilon}}({\mathbf{r}})$ one uses the localized shifting operators (64) to define a global shifting operator via integration:
[…]
$=\sum_{i}\big{\{}{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})\cdot\nabla_{i}-[\nabla_{i}{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})]:{\bf p}_{i}\nabla_{{\bf p}_{i}}\big{\}}.$ (81)
where the explicit form (81) is obtained from using $\boldsymbol{\sigma}({\mathbf{r}})$ according to Eq. (64) and carrying out the position integral over ${\mathbf{r}}$. The colon in Eq. (81) indicates a double tensor contraction, which equivalently is the trace of the product of the two matrices. The phase space shifting operator $\Sigma[{\boldsymbol{\epsilon}}]$ depends functionally on the shifting field ${\boldsymbol{\epsilon}}({\mathbf{r}})$ as we indicate by the brackets.
Figure 7: Pseudocodes that describe Monte Carlo in the standard form (Algorithm 1) and using phase space shifting (Algorithm 2). The standard Metropolis algorithm is based on trial displacement from ${\mathbf{r}}_{i},{\bf p}_{i}$ to ${\mathbf{R}}_{i},{\bf P}_{i}$, where the capitalization indicates the trial variables. In the shifted system the transformation is applied both to the current state, resulting in ${\mathbf{r}}^{\prime}_{i},{\bf p}^{\prime}_{i}$, as well as to the trial state, resulting in ${\mathbf{R}}_{i}^{\prime},{\bf P}_{i}^{\prime}$. The Metropolis criterion is then evaluated with respect to the shifted variables and these then also form the basis for the sampling of observables. The illustrations of the shifted system only display configurational degrees of freedom with the momentum variables being treated accordingly. The pseudocode for histogram sampling in the shifted system only describes the configurational strategy with corresponding momentum sampling being required for obtaining, e.g., the one-body phase space distribution function $f({\mathbf{r}},{\bf p})$ shown in Fig. 6.
The commutator relationship for full phase space shifting is given by
$[\Sigma[{\boldsymbol{\epsilon}}_{1}],\Sigma[{\boldsymbol{\epsilon}}_{2}]]$ $=\Sigma[{\boldsymbol{\epsilon}}_{\Delta}].$ (82)
The difference shifting vector field ${\boldsymbol{\epsilon}}_{\Delta}({\mathbf{r}}_{i})$ is thereby obtained from the given forms of ${\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})$ and ${\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})$ via the relation (27), previously obtained for pure position shifting, which we reproduce for convenience: ${\boldsymbol{\epsilon}}_{\Delta}({\mathbf{r}}_{i})={\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})\cdot[\nabla_{i}{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})]-{\boldsymbol{\epsilon}}_{2}({\mathbf{r}}_{i})\cdot[\nabla_{i}{\boldsymbol{\epsilon}}_{1}({\mathbf{r}}_{i})]$.
The commutator relation (82) forms a non-commutative Lie algebra due to the following three features: i) anti-symmetry, which originates from the very definition of the commutator, ii) bilinearity, which is induced by the linearity of the differential operator (81), and iii) the Jacobi identity: $[\Sigma_{1},[\Sigma_{2},\Sigma_{3}]]+[\Sigma_{2},[\Sigma_{3},\Sigma_{1}]]+[\Sigma_{3},[\Sigma_{1},\Sigma_{2}]]=0$, which can be shown by direct manipulation using Eq. (81). The functional arguments are suppressed in the notation such that $\Sigma_{1}=\Sigma[{\boldsymbol{\epsilon}}_{1}]$, $\Sigma_{2}=\Sigma[{\boldsymbol{\epsilon}}_{2}]$, and $\Sigma_{3}=\Sigma[{\boldsymbol{\epsilon}}_{3}]$.
By construction the shifting operators perform the following displacement to lowest order in the shifting field and its gradient:
$f({\mathbf{r}}^{\prime N},{\bf p}^{\prime N})$ $=f({\mathbf{r}}^{N},{\bf p}^{N})+\Sigma[{\boldsymbol{\epsilon}}]f({\mathbf{r}}^{N},{\bf p}^{N}),$ (83)
where we recall the relationship of original phase space variables ${\mathbf{r}}^{N},{\bf p}^{N}$ and their displaced versions ${\mathbf{r}}^{\prime N}$ and ${\bf p}^{\prime N}$ via the joint particle-resolved transformation (10) and (11).
For completeness, we can restore the localized shifting operators $\boldsymbol{\sigma}({\mathbf{r}})$ from $\Sigma[{\boldsymbol{\epsilon}}]$ by functional differentiation with respect to the shifting field according to:
$\boldsymbol{\sigma}({\mathbf{r}})$ $=\frac{\delta\Sigma[{\boldsymbol{\epsilon}}]}{\delta{\boldsymbol{\epsilon}}({\mathbf{r}})}.$ (84)
As a consequence, the following relationships hold between single and successive application of localized shifting and functional derivatives with respect to the shifting field:
$\boldsymbol{\sigma}({\mathbf{r}})f({\mathbf{r}}^{N},{\bf p}^{N})$ $=\frac{\delta f({\mathbf{r}}^{\prime N},{\bf p}^{\prime N})}{\delta{\boldsymbol{\epsilon}}({\mathbf{r}})}\Big{|}_{{\boldsymbol{\epsilon}}=0},$ (85)
$\boldsymbol{\sigma}({\mathbf{r}})\boldsymbol{\sigma}({\mathbf{r}}^{\prime})f({\mathbf{r}}^{N},{\bf p}^{N})$ $=\frac{\delta^{2}f({\mathbf{r}}^{\prime N},{\bf p}^{\prime N})}{\delta{\boldsymbol{\epsilon}}({\mathbf{r}})\delta{\boldsymbol{\epsilon}}({\mathbf{r}}^{\prime})}\Big{|}_{{\boldsymbol{\epsilon}}=0}$
$\quad+[\nabla\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})]\boldsymbol{\sigma}({\mathbf{r}})f({\mathbf{r}}^{N},{\bf p}^{N}).$ (86)
We have hence demonstrated that much of the favourable structure of coordinate shifting, as described in Sec. IV.3, is not only retained, see e.g. the corresponding operator algebra for phase space shifting (69) and coordinate shifting (53), but even formally simplified, see the simple self-adjoint nature of phase space shifting (68), as compared to the (slightly) more complex version (52) for coordinate shifting.
The mathematical structure that is encoded in the behaviour of differential transformations allows one to derive exact sum rules, as we have demonstrated. That the gauge invariance against phase space shifting is not a mere formal device but an intrinsic property of the statistical mechanics of many-body systems is arguably most strikingly demonstrated for finite shifting, as we turn to next.
V Computer Simulations
Finite phase space shifting is expressed by the general coordinate and momentum map (10) and (11) described in Sec. III.2. We wish to demonstrate that phase space averages are in practice invariant under this gauge transformation. As a prototypical example we consider hard rod particles of size $a$ in one spatial dimension. The hard core interaction assigns vanishing statistical weight to any configuration with two particles $i$ and $j$ having spatial distance $|x_{i}-x_{j}|<a$, where $x_{i}$ and $x_{j}$ are one-dimensional position coordinates. We refer the reader to Ref. sammueller2023whyNeural for a description of the emerging collective physics as described on the basis of classical density functional theory and using neural excess free energy functionals. We deem the hard rod model to constitute a fundamental test case. The results presented in Ref. mueller2024gauge for soft interparticle and external interaction potentials demonstrate the generality of our conclusions.
We use Monte Carlo simulations and choose two representative observables. One is the density “operator”, $\hat{A}=\hat{\rho}(x)=\sum_{i}\delta(x-x_{i})$ where $x$ denotes a generic position and we have cast the general definition (42) of $\hat{\rho}({\mathbf{r}})$ in one-dimensional form. We furthermore address the one-body phase space distribution function $f(x,p)=\langle\hat{f}(x,p)\rangle$, where the observable under consideration is $\hat{A}=\hat{f}(x,p)=\sum_{i}\delta(x-x_{i})\delta(p-p_{i})$, with $p$ being a generic momentum variable.
To carry out the simulations in the displaced system we use a one-dimensional shifting field given by the following specific form:
$\epsilon(x)=\epsilon_{0}\sin(4\pi x/L),$ (87)
where $\epsilon_{0}$ is an amplitude that governs the overall magnitude of the displacement, $L$ denotes the system size, which we choose as $L=10a$. The system is confined between two hard walls. Setting $\epsilon_{0}=0$ leads to vanishing shift and constitutes our baseline. We compare the corresponding results with those obtained for two different choices of values for $\epsilon_{0}$. We first set $\epsilon_{0}/a=0.5$, which is a safe choice below the threshold of invertibility and the resulting phase space transformation is bijective. As a further choice we take $\epsilon_{0}/a=1.5$, which is above the invertibility threshold, which we obtain in the present simple geometry as $\epsilon_{0}=L/(4\pi)=0.796$ from the requirement $d(x+\epsilon(x))/dx>0$.
The simulation results shown in Fig. 6 indicate that sampling in the validly displaced system $\epsilon_{0}/a=0.5$ (middle column) generates results that are numerically identical to those of the original system with $\epsilon_{0}/a=0$ (left column). The invalid transformation, upon using identical code and merely changing the amplitude of the shifting field to the value of $\epsilon_{0}/a=1.5$, yields corrupted results (right column). Such behaviour is expected for this case of a transform that is not bijective.
Pseudocode for the corresponding Metropolis Monte Carlo algorithm based on single particle moves is described in Fig. 7. We lay out the standard procedure as a reference (algorithm 1), together with the changes that implement phase space shifting both for the generation of the microstates and for the data acquisition via histogram filling (algorithm 2).
VI Conclusions
In conclusion we have described background and details for the gauge transformation of statistical mechanical microstates of Ref. mueller2024gauge . Thereby the thermal invariance against phase space shifting implies exact sum rules, as previously identified in a variety of different settings hermann2021noether ; hermann2022topicalReview ; hermann2022variance ; hermann2022quantum ; sammueller2023whatIsLiquid ; hermann2023whatIsLiquid ; robitschko2024any ; tschopp2022forceDFT . These include inhomogeneous thermal quantum systems hermann2022quantum , as well as classical liquids and more general soft matter states sammueller2023whatIsLiquid ; hermann2023whatIsLiquid . Previous work was based on variational methods and using invariance against the specific form of the displacement field ${\boldsymbol{\epsilon}}({\mathbf{r}})$ that parameterizes the phase space transformation. Here we have given background and details about the structure of the underlying invariance group including the algebra of phase space differential operators identified in Ref. mueller2024gauge . The relevant phase space operators represent infinitesimal versions of finite gauge transformations and the analysis of their properties allows one to reveal rich mathematical structure.
The gauge transformation affects particle position (10) and momentum (11) degrees of freedom. The vector field ${\boldsymbol{\epsilon}}({\mathbf{r}})$ that parameterizes the transform needs to be smooth and of a form that allows the transformation to be bijective. Specifically we require that the function ${\mathbf{r}}_{i}+{\boldsymbol{\epsilon}}({\mathbf{r}}_{i})$ is a diffeomorphism. With the position transform being very general, the momentum transform ensures that the differential phase space volume element is conserved and that the joint transformation (10) and (11) is canonical in the sense of classical mechanics goldstein2002 .
For finite displacement we have described in detail the geometric nature of position shifting. This transformation can be inverted and chained and hence establishes a group structure. Notwithstanding the apparent geometric simplicity, the group is non-commutative with the order of two consecutive shifts being relevant. Despite its more complex matrix multiplication structure, the corresponding momentum transform (11) remains entirely compatible with the mathematical structure of the position transform.
When considering the infinitesimal version of shifting much additional structure is revealed by formalizing relationships for differential operators that perform the shifting. The transformations are functionally dependent on the vector field ${\boldsymbol{\epsilon}}({\mathbf{r}})$ that parameterizes the transform. Functional differentiation yields corresponding commutator relationships for position-localized differential operators that act on the phase space of a many-body system. We have laid out the properties of pure position shifting and have compared against those of full phase space gauge transformation. Both versions of corresponding localized differential operators can be used as a basis for deriving exact statistical mechanical sum rules, as we have demonstrated. The mathematical derivation of these correlation function identities (sum rules) is greatly simplified by the use of the differential operator formalism.
The classical gauge theory that we have presented carries strong similarities with the theory of Lie groups and associated Lie algebras. In particular, the phase space shifting operators $\Sigma[{\boldsymbol{\epsilon}}]$ given by Eq. (81) form a non-commutative Lie algebra, as described by the commutator relation (82). For pure configuration shifting the corresponding Lie algebra is given by Eq. (29). The dependence on the shifting field ${\boldsymbol{\epsilon}}({\mathbf{r}})$ remains thereby explicit and it acts to identify the elements of the algebra.
In the present statistical mechanical setting it is useful to go further and eliminate the dependence on ${\boldsymbol{\epsilon}}({\mathbf{r}})$. We recall that ${\boldsymbol{\epsilon}}({\mathbf{r}})$ possesses the status of a mere gauge function, which does not affect the actual physics; see the similarities with gauge invariance in electrodynamics as described in Sec. II. Functional differentiation of the Lie algebra with respect to the shifting field constitutes a well-characterized route to eliminate the dependence on ${\boldsymbol{\epsilon}}({\mathbf{r}})$, which in turn generates spatial localization via Dirac distributions. The resulting operator commutator identities, see Eq. (69) for full phase space shifting and Eq. (34) for the purely configurational version, retain much of the favourable structure of a Lie algebra, see the occurrence of structure constants Eq. (35). However, the distribution character reaches beyond elementary Lie theory robbin2022book and it would be interesting in future work to explore connections with more abstract Lie concepts in modern mathematical treatments of the subject.
We have described several implications of the gauge invariance for simulation methods focusing on Monte Carlo sampling. Choosing a specific form of the shifting field and displacing the particles in a prototypical confined one-dimensional hard core system allows one to give a standalone demonstration of gauge invariance. Applying the transformation (10) and (11) goes in practice beyond a mere change of coordinates, as this procedure genuinely alters the resulting Markov chain due to the evaluation of acceptance probabilities in the virtually displaced system. Nevertheless, the invariance property ensures that thermal averages remain valid despite the seemingly violated requirement of detailed balance. We recall that acceptance rates differ in the displaced and original systems, which serves as a practical indicator for the differences in sampled states upon, as theoretically predicted, retaining identical one-body density profile and phase space distribution function, which we took as representative of the behaviour of general observables.
In future work, it would be interesting to investigate the potential use of gauge invariance in machine learning of neural density functionals sammueller2023neural ; sammueller2023whyNeural , for constructing analytical density functional approximations gul2024testParticle , cross fertilization with hyperdensity functional theory sammueller2024hyperDFT ; sammueller2024whyhyperDFT , construction of advanced simulation schemes borgis2013 ; delasheras2018forceSampling ; coles2019 ; coles2021 ; rotenberg2020 ; purohit2019 ; schultz2016 ; trokhymchuk2019 ; schultz2019 ; purohit2019 , in particular based on mapped averaging moustafa2015 ; schultz2016 ; moustafa2017jctp ; moustafa2017prb ; lin2018 ; moustafa2019 ; schultz2018 ; purohit2018 ; schultz2019 ; purohit2019 ; trokhymchuk2019 ; purohit2020 ; moustafa2022 , and application to mixtures matthes2024mix .
Acknowledgements.
We thank Sophie Hermann for useful discussions. This work is supported by the DFG (Deutsche Forschungsgemeinschaft) under project 551294732.
Appendix A Functional derivatives in electrodynamics
We denote the transformed action, as given via the right hand side of Eq. (8), by $S_{\rm ext}[\varphi]$. The brackets indicate functional dependence, which arises as the value of the action depends a priori on the specific form of the gauge function $\varphi(x)$. We can express the gauge invariance as the identity
$S_{\rm ext}=S_{\rm ext}[\varphi],$ (88)
where the left hand side denotes the original action (5), which is clearly independent of $\varphi(x)$. The invariance equation (88) holds irrespective of the form of the gauge function $\varphi(x)$. Hence a valid identity is retained upon functionally differentiating both sides by $\varphi(x)$. The left hand side of Eq. (88) then vanishes as there is no dependence on the gauge function.
[…]
While requiring slightly more steps than the the route considered in the main text, the present method via functional calculus is arguably as generally applicable as is ordinary multivariate calculus and hence there are no principal limits in terms of complexity of the functional dependence under investigation. Moreover, as was demonstrated recently for statistical functionals sammueller2023neural ; sammueller2023whyNeural ; stierle2024autodiff , powerful software tools that implement automatic differentiation baydin2018autodiff can be used.
For completeness, the free field action (4) is unaffected by the gauge transformation, as $F_{\lambda\nu}(x)$ is already an invariant. We demonstrate this explicitly as follows. The free field action (4) acquires the following apparent change under the gauge transformation (3) according to
[…]
From invariance of the free field action one can conclude that $\partial_{\lambda}\partial_{\nu}F^{\lambda\nu}(x)=0$, which already follows in an elementary way from the antisymmetry $F^{\lambda\nu}(x)=-F^{\nu\lambda}(x)$.
Again an alternative derivation can be based on functional differentiation which, following steps analogous to those described above for the external action, gives the same result: $\delta S_{\rm free}[\varphi]/\delta\varphi(x)=\partial_{\lambda}\partial_{\nu}F^{\lambda\nu}(x)/\mu_{0}=0$.
To make the covariant formulation more explicit, the relationship of $F_{\nu\lambda}(x)$ to the electrical field and to the magnetic induction is via the explicit components: $F_{0\lambda}(x)=(0,E_{x}/c,E_{y}/c,E_{z}/c)$, $F_{1\lambda}(x)=(-E_{x}/c,0,-B_{z},B_{y})$, $F_{2\lambda}(x)=(-E_{y}/c,B_{z},0,-B_{x})$ $F_{3\lambda}(x)=(-E_{z}/c,-B_{y},B_{x},0)$, where ${\bf E}({\mathbf{r}},t)=(E_{x},E_{y},E_{z})$ and ${\bf B}({\mathbf{r}},t)=(B_{x},B_{y},B_{z})$. For completeness, the entire field tensor reads as follows:
$F_{\nu\lambda}(x)$ $=\left({\begin{array}[]{cccc}0&E_{x}/c&E_{y}/c&E_{z}/c\\ -E_{x}/c&0&-B_{z}&B_{y}\\ -E_{y}/c&B_{z}&0&-B_{x}\\ -E_{z}/c&-B_{y}&B_{x}&0\\ \end{array}}\right).$ (94)